OpenStudy (anonymous):

The sum of the roots of the equation 1/2x^2-5/4x-3=0 is: 5/2 2 1/2

OpenStudy (anonymous):

5/4 isn't one of the answer choices.

OpenStudy (anonymous):

you are after the sum of the roots and not the roots themselves. Is it?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

no need to find the roots then add them up

OpenStudy (anonymous):

You're not making any sense.

OpenStudy (anonymous):

it should be: in your case it is: (5/4)/(1/2)=5/2

OpenStudy (anonymous):

it is: -b/a

OpenStudy (anonymous):

Ahh okay... You put something different the first time lol

OpenStudy (anonymous):

yup. am now correct :)

OpenStudy (anonymous):

Okay, so what about the product for the same equation?

OpenStudy (anonymous):

i was surprised by your beauty

OpenStudy (anonymous):

Lol well thank you. Will you help me some more please?

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

product is -6

OpenStudy (anonymous):

what else?

OpenStudy (anonymous):

I PM'd you :-)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ask here

OpenStudy (anonymous):

Which of the following statements are true of the discriminant? The discriminant is b^2 – 4ac in the quadratic formula. The discriminant is the denominator in the quadratic formula. The discriminant is the radicand of the quadratic formula. The discriminant is the value under the radical in the quadratic formula. The discriminant is the value in front of the radical in the quadratic formula.

OpenStudy (anonymous):

first

OpenStudy (anonymous):

I know it's the first and third, but I wanted to see if there were more.

OpenStudy (anonymous):

fourth also

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. x^2 - 4x + 4 = 0 double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

double root

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. x^2 - 5x + 7 = 0 double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

imaginary

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. x^2 - 5x - 4 = 0 double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

real and irrational

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 16x^2 + 8x + 1 = 0 double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

double root

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. x2 + 9x + 14 = 0 double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

imaginary

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 3x2 - 10 = 0 double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

sorry. the one before is: real rational

OpenStudy (anonymous):

last: real irrational

OpenStudy (anonymous):

medals :)

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 2 = x2 + 5x double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

real irrational

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 2m2 + 3 = m double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

imaginary

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 6x2 + 13x + 6 = 0 double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

real rational

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 2x2 + 7x + 6 = 0 double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

real rational

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 7x2 + 1 = 5x double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

imaginary

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 7x2 + 3 = 8x double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

do u still have many

OpenStudy (anonymous):

it is boring :)

OpenStudy (anonymous):

4 left

OpenStudy (anonymous):

ok cutie

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 7x2 + 3 = 8x double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

real irrational

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. 3x2 - 2 + 7x = 0 double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

ops. the one before is imaginary

OpenStudy (anonymous):

last real irrational

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. x2 - 6x + 9 = 0 double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

double

OpenStudy (anonymous):

Use the discriminant to describe the roots of each equation. Then select the best description. x2 - 6x + 12 = 0 double root real and rational root real and irrational root imaginary root

OpenStudy (anonymous):

imaginary

OpenStudy (anonymous):

Will you please please please help me with another assignment? :)

OpenStudy (anonymous):

ops. not now

OpenStudy (anonymous):

Pleaseee. I'm begging you

OpenStudy (anonymous):

it is my pleasure :)

OpenStudy (anonymous):

So you will help?

OpenStudy (anonymous):

i see no medals

OpenStudy (anonymous):

I will post a new thread.